Suppose all licence plates in a certain region have three letters and three digits. If a plate is chosen at random. What is the probability that all three letters and all the numbers on the plate Will be different?
1
Expert's answer
2020-07-02T19:05:44-0400

Solution.

Suppose there is "n" letters and "k" digits ("n \geq 3, k \geq 3)", that can be on licence plate. And there are no any special rules, such as "digits 000 is forbidden".

Amount of all possible plates is "n^3\cdot k^3."

Amount of plates with unique letters and digits is "n(n-1)(n-2)\cdot k(k-1)(k-2)."

So the probability is "p = \dfrac{n(n-1)(n-2)\cdot k(k-1)(k-2)}{n^3\cdot k^3}=\dfrac{(n-1)(n-2)\cdot(k-1)(k-2)}{n^2\cdot k^2} \\[0.1cm]"

If "n = 26" and "k = 10", then "p=\dfrac{25\cdot24\cdot9\cdot8}{26^2\cdot10^2}\approx 0.639."


Answer.

0.639